OpenStudy (anonymous):

Simplify the following expression. x − 4y + 2z − 3x − 3y + 2z = ?

OpenStudy (anonymous):

Just combine like terms

OpenStudy (anonymous):

ok

OpenStudy (nathan917):

The solution to this is simple, combine like terms like this: 4x-3x= x -3y-y= -4y 2z+3z= 5z There for the answer is x-4y+5z @mattdagrinch12

OpenStudy (anonymous):

That's not right nathan, I don't know where you got those values, but those are not the values from the problem

OpenStudy (nathan917):

@mattdagrinch12 @bfpGunz -5y+5z -5(y -z) Did the question rong srry.